OpenStudy (anonymous):

HELP ME PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Solve the equation for y: 12x – 4y = 8. y = –2 + 3x y = 2 – 3x y = –8 + 12x y = 12 – 12x

OpenStudy (mertsj):

Subtract 12x from both sides.

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

nothing..... im confused....

OpenStudy (mertsj):

Do you know what the two sides of an equation are?

OpenStudy (anonymous):

nope....

OpenStudy (zzr0ck3r):

great question:)

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

The equal sign divides the sides of the equation. In your equation the left side is 12x-4y. What is the right sides?

OpenStudy (anonymous):

razzy if you have 12x - 12x...what do you get?

OpenStudy (mertsj):

side

OpenStudy (mertsj):

Ok. Musical rose wants to help you now. Good bye.

OpenStudy (anonymous):

its -4y=8? rite?

OpenStudy (zzr0ck3r):

12x - 4y = 8 we want Y on one side and the rest on the other if we subtract 12x from both sides 12x-12x-4y=8-12x -4y=8-12x then to get y alone we need to devide by the number next to y (-4) (-4/-4)y = (8/-4)-(12/-4)x y = -2 + 3x does this make since?

OpenStudy (zzr0ck3r):

add/subtract terms away from the side with y untill you get one term with y on one side, then devide away the rest of the term to get y by itself

OpenStudy (anonymous):

ok wait im confused on whats the answer......

OpenStudy (zzr0ck3r):

you need to worry more about how to get there not about what the answer is. If you understand what I did it should be obviouse what the answer is in my explination

OpenStudy (anonymous):

oh then well i dont understand what u did

OpenStudy (zzr0ck3r):

if you have y+3 = 7 what does y equal?

OpenStudy (zzr0ck3r):

ok read what I said and tell me what part you dont udnerstand

OpenStudy (zzr0ck3r):

line by line